5how that the vecters uvd besis W forn ay2 lay2 the vector t ih respect bWrite to the basis += in
Answer #1

9. show that the vectors uvfW foems a basis. a) u = (3) are v-(6),w=(7) =) Let s= {U, V W} then sis basis for IRS iff i s spa20-3btza 15 =) |K3 = 26-36+39 K2 - 2 kg = bna ke = bac + 2 kg = bag +8 (20-3b +39) = 5(b-a) + 2(-36t3a - 5b-5a721-36+3 a K2=(3,4,0) +(0, 2X2,9x2)+(23,6283,333) = 69,0,0). & t3 = 0 & +222222=0 342 +343=0. now, 21 +23=0 =) & =-23 0 & 322+ 323=0 - L2=-